OpenStudy (mitu12):

There are 13 animals in the bar. Some are chickens and some are pigs. There are 40 legs in all. How many of each animal are there?

OpenStudy (anonymous):

they really stretch for these word problems, don't they? try \[x+y=13\\ 2x+4y=40\]

OpenStudy (anonymous):

You can make 2 equations based off this equation. A chicken has two legs and a pig has four. x+y=13 2x+4y=40 Solve these two systems of equations x=13-y 2(13-y)+4y=40 26-2y+4y=40 2y+26+40 Subtract 26 on both sides 2y=14 Divide by 2 on both sides y=7 x+7=13 x=6 So there are 6 chickens and 7 pigs.

OpenStudy (mitu12):

thanks